# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ 1 Introduction
Identifying the nature and properties of dark matter remains a crucial open problem in fundamental physics, with major implications for astrophysics and for the evolution of the universe at the cosmological level [1]. By now, the hypothesis of the existence of dark matter is supported by observational evidence from many effects, on widely separated scales, ranging from the motion, structure, and evolution of many galaxies, to weak-gravitational-lensing effects in galaxy clusters, to the cosmic microwave background, to other effects relevant for cosmology [2, 3, 4, 5, 6, 7]. In particular, Big-Bang nucleosynthesis and observations of the cosmic microwave background indicate that roughly a quarter of the present energy density of the universe is accounted for by dark matter; in addition, the latter also plays a crucial rôle in explaining existence of inhomogeneities associated with the existence of structures like galaxies and clusters: if a matter species decoupled from photons much before baryons did during the evolution of the early universe, it would have allowed density perturbations to grow sufficiently, before the baryon recombination time. Even though some of these observations may be explained by modified theories of gravity [8], others cannot [9, 10, 11, 12], leaving a particle interpretation of dark matter as the most likely option [13, 14, 15].
@@
Among many models that have been proposed (for an overview, see, for example, ref. [25]), the possibility that dark matter is made of composite states of a novel, strongly coupled theory is especially appealing, both for theoretical and for phenomenological reasons [26]. Even at a very heuristic level, one should note that already the largest fraction of visible mass in the universe is generated dynamically by a strongly coupled gauge theory, namely quantum chromodynamics (QCD), so the same mechanism may be at work for dark matter, too—and, much like the dynamically generated hadronic scale in quantum chromodynamics, the mass scale of these composite states would not be affected by the naturalness problem. Such composite states are “dark” if they are globally neutral under SM, even if their elementary constituents had non-vanishing SM charges. Moreover, their stability against decays could be due to some internal symmetry: if the leading operators that violate the latter have a large dimension, then the induced decay rate can be sufficiently small to make the lifetime of dark matter comparable with or longer than the age of the universe; note that this is analogous to the accidental baryon-number $\operatorname{\mathrm{U}}(1)$ symmetry of the renormalizable SM Lagrangian, which, being broken only by dimension-$6$ operators, explains the very high lower bound on the proton lifetime [27]. In addition, this mechanism could also explain the (so far) experimentally unobserved interaction between dark and ordinary matter in terms of the suppression of the corresponding interaction terms by inverse powers of the dynamically generated dark confinement scale. A dark confining gauge theory supporting bound states with sufficiently large self-interaction cross section, but negligible annihilation or dissipation, may be consistent with observational constraints related to structure formation. Finally, in the literature it has been suggested that an interpretation of dark matter in terms of a theory akin to QCD could even explain the fact that the contributions to the total energy density of the universe from visible matter and from dark matter are of comparable magnitude [28].
@@
A particularly simple realization of this type of dark-matter models is given by a confining non-Abelian gauge theory, without matter fields—a model similar to the purely gluonic sector of QCD. While the fundamental degrees of freedom of this model correspond to massless, “gluon-like” particles not coupled to the Standard Model (SM) fields, the confinement phenomenon and the dynamical generation of a mass gap implies that the physical states in the spectrum of the theory are massive “glueball” states, which are neutral under the fundamental “color-like” charge of the theory and under the SM. In particular, the lightest such state, which has spin, parity and charge quantum numbers $J^{PC}=0^{++}$, can be a dark matter candidate [29, 30], as it is predicted to be completely neutral under the SM charges, to be stable on cosmological timescales by virtue of the confining nature of the new gauge interaction, and to interact with the SM particles only gravitationally, thereby providing a very secluded type of dark matter. Among the phenomenologically appealing features of this type of models, it is worth remarking that they naturally lead to a type of dark matter that could simultaneously be warm and self-interacting, that would account for the lack of direct experimental detection at colliders, and that features a dynamical mechanism for mass generation free from the naturalness problem. In addition, the model depends only on a very limited number of free parameters and, therefore, has strong predictive power: once the gauge group is specified, all dimensionful physical quantities depend on a single energy scale. As we will discuss in detail below, this means, for example, that the masses of the entire spectrum of the theory are proportional to the lightest one, through ratios that are fixed and exactly computable from the microscopic Lagrangian of the theory – as long as a non-perturbative calculation tool is available. In turn, the running of the renormalized coupling of the theory (in a given scheme) is specified by a $\Lambda$ parameter, which is also proportional to the mass of the lightest physical state in the spectrum; similarly, the force between a pair of probe sources in the fundamental representation of the gauge group at asymptotically large distances is proportional to the square of the mass of the lightest “glueball” of the theory.
@@
Further interesting implications of a glueball dark-matter model can be derived, once one considers its behavior at a finite temperature $T$: in particular, at a sufficiently high temperatures the color-singlet states cease to exist and give way to a plasma of gluon-like quasi-particle excitations interacting with each other through thermally screened long-distance forces. Again, the value of the critical deconfinement temperature $T_{\mbox{\tiny{c}}}$ separating these two phases is proportional to the mass gap of the theory, and, if the confining and the plasma phase are separated by a first-order transition, the associated latent heat per unit volume is proportional to the fourth power of the mass gap. Remarkably, the liberation of a large number of light degrees of freedom at high temperatures would have an impact on the total density of free energy and other thermodynamic quantities in the early universe, affecting its cosmological cooling rate.
@@
The simplest formulation of a glueball-like model for dark matter is one in which not only the “dark glueballs” but also their constituent “dark gluons” are neutral under the SM. However, it is interesting to note that, if instead the constituents do carry SM charges, when they are deconfined at the very high temperatures of the early universe they could interact sufficiently strongly with the SM particles: in principle, this could explain the fact that the present densities of baryonic matter and dark matter are of the same order of magnitude, and may also have interesting implications for the baryon asymmetry of the universe [31, 32, 33].
@@
In the present work, we study the viability as a dark-matter model of a Yang-Mills theory based on local invariance under the $\operatorname{\mathrm{Sp}}(2)$ compact symplectic group, which is the compact real form of the symplectic group over the field of complex numbers $\operatorname{\mathrm{Sp}}(4,\operatorname{\mathbb{C}})$, and the intersection of the latter with the unitary group $\operatorname{\mathrm{U}}(4)$. [^1] Note that the compact symplectic groups like the one considered in this work are sometimes denoted as $\operatorname{\mathrm{Sp}}(2N)$, or, alternatively, as $\operatorname{\mathrm{USp}}(2N)$, in order to emphasize that, in the defining representation, they are given by matrices of size $2N\times 2N$. In the present work, we use, instead, the convention in which they are denoted as $\operatorname{\mathrm{Sp}}(N)$, as was done, for example, in ref. [34]. In particular, we focus on the behavior of this model at temperatures of the order of and above the deconfinement transition, and determine its equation of state non-perturbatively, by means of numerical calculations in the lattice regularization [35].
@@
An interesting feature of the model that we consider, is that the dark sector may have undergone a strong first-order thermal phase transition in the early universe, possibly producing gravitational waves [36, 37, 38, 39]. In addition to the gravitational waves produced by non-equilibrium processes at a discontinuous transition, it should be remarked that any plasma in thermal equilibrium produces gravitational waves, with an emission rate proportional to the shear viscosity [40]; while the latter can become relatively large at weak coupling, this contribution to the gravitational-wave spectrum remains subdominant with respect to those generated at a first-order phase transition. Gravitational waves produced in a dark sector contribute to the total energy density (and to the effective number of massless degrees of freedom), so the experimental detection of gravitational waves with a characteristic spectrum could provide useful hints about new physics. For example, a first-order phase transition at temperatures of the order of hundreds $\mathrm{GeV}$, i.e., the scale characteristic of electroweak physics, could generate gravitational waves with peak frequencies of the order of $10^{-3}$ Hz, which are potentially within the reach of space-based gravitational-wave observatories [41, 42, 43, 44, 45, 46, 47]. Given that the SM predicts only a smooth crossover at the electroweak scale [48, 49], the experimental detection of gravitational waves in this frequency range could be interpreted as evidence of physics beyond the Standard Model.
@@
Models featuring dark matter candidates in the form of glueball states have been studied in several recent articles; a partial list of references includes refs. [63, 64, 65, 66, 67, 68, 69, 70, 62, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94]. This family of models can be thought of examples of the “hidden valley” paradigm [95, 96, 97] (which admit realizations in various theoretical scenarios [98, 99, 100, 101, 102, 103]) and offer phenomenologically attractive features; in particular, while they evade current bounds from colliders [104], they could be coupled to SM particles through portal interactions [105, 106, 107, 108].
The spectrum and related properties of glueballs in $\operatorname{\mathrm{SU}}(N)$ gauge theories have been studied in the lattice regularization in various works [109, 110, 111, 112, 113, 114, 115, 116, 117, 23, 118, 119], but the literature on analogous computations in $\operatorname{\mathrm{Sp}}(N)$ gauge theories is more limited. In fact, with the exception of the smallest symplectic group $\operatorname{\mathrm{Sp}}(1)$ (which is nothing but $\operatorname{\mathrm{SU}}(2)$, and gauge theories based on this group – with or without dynamical matter fields – have already been studied very extensively in lattice simulations, both for the qualitative similarities with QCD and for the potential relevance for physics beyond the SM: for a recent review, see ref. [120, subsection 5.2]), the first lattice study of symplectic gauge theories was reported in ref. [34]; during the past few years, research on this subject has gained significant momentum [121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139].
@@
The structure of this article is as follows. After discussing the formulation of $\operatorname{\mathrm{Sp}}(N)$ gauge theories in the lattice regularization and the method to compute the equation of state in section 2, we present our results for the equilibrium thermodynamic properties of $\operatorname{\mathrm{Sp}}(2)$ theory in section 3, while in section 4 we discuss the possible interpretation of these results in terms of a model for dark matter, and their implications, before recapitulating our findings in section 5. Some general properties of symplectic groups are discussed in appendix A, while appendix B is devoted to technical details about our Markov-chain Monte Carlo algorithm. Throughout this paper we assume natural units, whereby the speed of light in vacuum, the reduced Planck’s constant and Boltzmann’s constant are set to unity.
@@ 2 Formulation of the model
Our non-perturbative study of the equation of state of $\operatorname{\mathrm{Sp}}(2)$ Yang-Mills theory is carried out through a regularization of the theory on a Euclidean, isotropic, hypercubic lattice $\Lambda$ of spacing $a_{\mbox{\tiny{lat}}}$. The physical extent of the system in each of the three spatial directions (labeled by indices $1$, $2$ and $3$) is denoted as $L=a_{\mbox{\tiny{lat}}}N_{s}$, while the size of the Euclidean-time direction (labeled by the index 00) is denoted as $a_{\mbox{\tiny{lat}}}N_{\tau}$; furthermore, let $V=L^{3}$ denote the spatial volume of the system. Periodic boundary conditions are imposed along the direction of the four main axes, so that the physical temperature $T$ of the system is given by $T=1/(a_{\mbox{\tiny{lat}}}N_{\tau})$. The simulations at finite temperature are performed on lattices with $N_{s}/N_{\tau}\gtrsim 4$ to suppress finite-volume corrections, while $N_{\tau}=N_{s}$ for the simulations at $T\simeq 0$. For a theory based on a generic $\operatorname{\mathrm{Sp}}(N)$ gauge group, the fundamental degrees of freedom of the theory regularized on the lattice are a set of $4N_{s}^{3}N_{\tau}$ square matrices in the defining representation of the $\operatorname{\mathrm{Sp}}(N)$ group; these matrices have size $2N$ and are defined on the oriented bonds between nearest-neighboring lattice sites; in particular, we denote the matrix on the link from the site $x$ to the site $x+a_{\mbox{\tiny{lat}}}\vec{\mu}$ as $U_{\mu}(x)$. The Euclidean action of the lattice theory is taken to be the Wilson action [35]
@@
In order to express the results extracted from lattice simulations (which are dimensionless numbers, i.e., are measured in units of the appropriate power of the lattice spacing) as physical quantities, one has to set the scale, i.e. to establish the correspondence between the values of the bare coupling of the lattice theory $g_{0}$ (or, equivalently, the $\beta$ parameter) and the lattice spacing. In this work we set the scale non-perturbatively, by determining the critical $\beta$ at which the deconfinement transition occurs for different values of $N_{\tau}$, interpreting the corresponding values of $1/(a_{\mbox{\tiny{lat}}}N_{\tau})$ as the deconfinement temperature $T_{\mbox{\tiny{c}}}$, and interpolating the numerical results obtained from this procedure to express the lattice spacing as a continuous function of $\beta$ in the range of values that were simulated. [^2] Since the physics of purely gluonic non-Abelian gauge theories in four spacetime dimensions (and without a topological $\theta$ term) depends only on one parameter, setting the scale using a different dimensionful quantity would lead to the same $a_{\mbox{\tiny{lat}}}(\beta)$ function, up to small corrections due to discretization effects. In turn, the determination of the $\beta$ value corresponding to the critical deconfinement temperature is carried out by studying the fourth-order Binder cumulant [158]
@@
$\mathcal{L}\left(\vec{x}\right)=\frac{1}{4}\operatorname{\mathrm{Tr}}\prod_{k=0}^{N_{\tau}-1}U_{0}\left(x+ka_{\mbox{\tiny{lat}}}\hat{0}\right)$ (2.9)
@@
The critical $\beta$ value in the thermodynamic limit is determined through a finite-size scaling analysis, from the crossing of the curves $B_{4}(\beta)$ obtained at different values of the linear spatial size of the system $L$. An alternative quantity that can be used to identify the $\beta$ value corresponding to deconfinement is the susceptibility associated with the spatially-averaged Polyakov loop: [^3] Note that the definition of $\chi_{\mathcal{L}}$ is actually based on the absolute value of $\mathcal{L}$ (rather than $\mathcal{L}$ itself) because all simulations are carried out in finite volume and at finite lattice spacing, and thus involve a strictly finite number of degrees of freedom; as a consequence, any non-vanishing expectation value of $\mathcal{L}$ would simply be an artefact of the finiteness of the sample of configurations. This problem is solved by using $|\mathcal{L}|$, instead of $\mathcal{L}$, in the definition of $\chi_{\mathcal{L}}$.
@@
[…]
@@
[…]
@@
Note that, from the pressure and the energy density, it is also possible to derive the square of the speed of sound $c_{\mbox{\tiny{s}}}$:
@@
$L_{\mbox{\tiny{h}}}=\lim_{V\to\infty}\frac{6}{a_{\mbox{\tiny{lat}}}^{4}}\frac{\partial\beta}{\partial\ln a_{\mbox{\tiny{lat}}}}\left(\langle P\rangle_{T_{\mbox{\tiny{c}}}^{-}}-\langle P\rangle_{T_{\mbox{\tiny{c}}}^{+}}\right).\textasciidieresis$ (2.19)
@@
Finally, before presenting our original results for the thermodynamics of the $\operatorname{\mathrm{Sp}}(2)$ theory, we discuss the classification of different glueball states whose masses can be computed on the lattice; as will be discussed in section 3, this is relevant for our discussion of the equation of state of the theory in the confining phase, which can be modelled in terms of a gas of massive, and approximately non-interacting, glueballs. On a hypercubic lattice, the symmetry under spatial rotations is broken down to rotations by integer multiples of $\frac{\pi}{2}$ around the three main spatial axes, so that glueball states are actually classified according to the irreducible representations of the orientation-preserving octahedral group (the symmetry group of the cube, with $24$ elements falling into five distinct conjugacy classes), which is isomorphic to the group of permutations of four elements $S_{4}$. The latter has five conjugacy classes (respectively containing the identity, the six permutations that can be written as a single cycle of length two, the three permutations that can be written as the product of two disjoint cycles of length two, the eight permutations expressed by a cycle of length three, and the six permutations expressed by a cycle of length four) and hence five irreducible representations; Burnside’s theorem fixes their dimensions to be $1$, $1$, $2$, $3$, and $3$: they are the trivial representation, the sign representation, the two-dimensional irreducible representation, the standard representation, and the product of the standard and sign representations. When regarded as irreducible representations of the orientation-preserving octahedral group, they are usually denoted as $A_{1}$, $A_{2}$, $E$, $T_{1}$, and $T_{2}$, respectively. The restriction of irreducible representations of the $\operatorname{\mathrm{SO}}(3)$ group to the orientation-preserving octahedral group, i.e., the subduced representations, can be decomposed into tensor sums of the various irreducible representations $A_{1}$, $A_{2}$, $E$, $T_{1}$, and $T_{2}$ (for a discussion, see, e.g., refs. [162, 115]). In particular, this implies that the $12J+$ 1 states of a spin-$J$ representation in the continuum get split among different irreducible representations of the orientation-preserving octahedral group; for example:

### v2

# Thermal evolution of dark matter and gravitational-wave production in the early universe from a symplectic glueball model

@@ 1 Introduction
Identifying the nature and properties of dark matter remains a crucial open problem in fundamental physics, with major implications for astrophysics and for the evolution of the universe at the cosmological level [1]. By now, the hypothesis of the existence of dark matter is supported by observational evidence from many effects, on widely separated scales, ranging from the motion, structure, and evolution of many galaxies, to weak-gravitational-lensing effects in galaxy clusters, to the cosmic microwave background, to other effects relevant for cosmology [2, 3, 4, 5, 6, 7]. In particular, Big-Bang nucleosynthesis and observations of the cosmic microwave background indicate that roughly a quarter of the present energy density of the universe is accounted for by dark matter; in addition, the latter also plays a crucial rôle in explaining the inhomogeneities associated with the existence of structures like galaxies and clusters: if a matter species decoupled from photons much before baryons did during the evolution of the early universe, it would have allowed density perturbations to grow sufficiently, before the baryon recombination time. Even though some of these observations may be explained by modified theories of gravity [8], others cannot [9, 10, 11, 12], leaving a particle interpretation of dark matter as the most likely option [13, 14, 15].
@@
Among many models that have been proposed (for an overview, see, for example, ref. [25]), the possibility that dark matter is made of composite states of a novel, strongly coupled theory is especially appealing, both for theoretical and for phenomenological reasons [26]. Even at a very heuristic level, one should note that already the largest fraction of visible mass in the universe is generated dynamically by a strongly coupled gauge theory, namely quantum chromodynamics (QCD), so the same mechanism may be at work for dark matter, too—and, much like the dynamically generated hadronic scale in quantum chromodynamics, the mass scale of these composite states would not be affected by the naturalness problem. The composite states of a strongly coupled sector are “dark” if they are globally neutral under SM, even if their elementary constituents had non-vanishing SM charges. Moreover, their stability against decays could be due to some internal symmetry: if the leading operators that violate the latter have a large dimension, then the induced decay rate can be sufficiently small to make the lifetime of dark matter comparable with or longer than the age of the universe; note that this is analogous to the accidental baryon-number $\operatorname{\mathrm{U}}(1)$ symmetry of the renormalizable SM Lagrangian, which, being broken only by dimension-$6$ operators, explains the very high lower bound on the proton lifetime [27]. In addition, this mechanism could also explain the (so far) experimentally unobserved interaction between dark and ordinary matter in terms of the suppression of the corresponding interaction terms by inverse powers of the dynamically generated dark confinement scale. A dark confining gauge theory supporting bound states with sufficiently large self-interaction cross section, but negligible annihilation or dissipation, may be consistent with observational constraints related to structure formation. Finally, in the literature it has been suggested that an interpretation of dark matter in terms of a theory akin to QCD could even explain the fact that the contributions to the total energy density of the universe from visible matter and from dark matter are of comparable magnitude [28].
@@
A particularly simple realization of this type of dark-matter models is given by a confining non-Abelian gauge theory, without matter fields—a model similar to the purely gluonic sector of QCD. While the fundamental degrees of freedom of this model correspond to massless, “gluon-like” particles not coupled to the Standard Model (SM) fields, the confinement phenomenon and the dynamical generation of a mass gap implies that the physical states in the spectrum of the theory are massive “glueball” states, which are neutral under the fundamental “color-like” charge of the theory and under the SM. In particular, the lightest such state, which has spin, parity and charge quantum numbers $J^{PC}=0^{++}$, can be a dark matter candidate [29, 30], as it is predicted to be completely neutral under the SM charges, to be stable on cosmological timescales by virtue of the confining nature of the new gauge interaction, and to interact with the SM particles only gravitationally, thereby providing a very secluded type of dark matter. Among the phenomenologically appealing features of this type of models, it is worth remarking that they naturally lead to a type of dark matter that could simultaneously be warm and self-interacting, and that would not be ruled out by the absence of direct experimental detection at colliders. In addition, the model depends only on a very limited number of free parameters and, therefore, has strong predictive power: once the gauge group is specified, all dimensionful physical quantities depend on a single energy scale. As we will discuss in detail below, this means, for example, that the masses of the entire spectrum of the theory are proportional to the lightest one, through ratios that are fixed and exactly computable from the microscopic Lagrangian of the theory – as long as a non-perturbative calculation tool is available. In turn, the running of the renormalized coupling of the theory (in a given scheme) is specified by a $\Lambda$ parameter, which is also proportional to the mass of the lightest physical state in the spectrum; similarly, the force between a pair of probe sources in the fundamental representation of the gauge group at asymptotically large distances is proportional to the square of the mass of the lightest “glueball” of the theory.
@@
Further interesting implications of a glueball dark-matter model can be derived, once one considers its behavior at a finite temperature $T$: in particular, at sufficiently high temperatures the color-singlet states cease to exist and give way to a plasma of gluon-like excitations interacting with each other through thermally screened long-distance forces. Again, the value of the critical deconfinement temperature $T_{\mbox{\tiny{c}}}$ separating these two phases is proportional to the mass gap of the theory, and, if the confining and the plasma phase are separated by a first-order transition, the associated latent heat per unit volume is proportional to the fourth power of the mass gap. Remarkably, the liberation of a large number of light degrees of freedom at high temperatures would have an impact on the total density of free energy and other thermodynamic quantities in the early universe, affecting its cosmological cooling rate.
@@
The simplest formulation of a glueball-like model for dark matter is one in which not only the “dark glueballs” but also their constituent “dark gluons” are neutral under the SM. However, it is interesting to note that, if instead the constituents do carry SM charges, when they are deconfined at the very high temperatures of the early universe they could interact sufficiently strongly with the SM particles: in principle, this could explain the fact that the present densities of baryonic matter and dark matter are of the same order of magnitude, and may also have interesting implications for the baryon asymmetry of the universe [31, 32, 33, 34].
@@
In the present work, we study the viability as a dark-matter model of a Yang-Mills theory based on local invariance under the $\operatorname{\mathrm{Sp}}(2)$ compact symplectic group, which is the compact real form of the symplectic group over the field of complex numbers $\operatorname{\mathrm{Sp}}(4,\operatorname{\mathbb{C}})$, and the intersection of the latter with the unitary group $\operatorname{\mathrm{U}}(4)$. [^1] Note that the compact symplectic groups like the one considered in this work are sometimes denoted as $\operatorname{\mathrm{Sp}}(2N)$, or, alternatively, as $\operatorname{\mathrm{USp}}(2N)$, in order to emphasize that, in the defining representation, they are given by matrices of size $2N\times 2N$. In this work, we use, instead, the convention in which they are denoted as $\operatorname{\mathrm{Sp}}(N)$, as was done, for example, in ref. [35]. In particular, we focus on the behavior of this model at temperatures of the order of and above the deconfinement transition, and determine its equation of state non-perturbatively, by means of numerical calculations in the lattice regularization [36].
@@
An interesting feature of the model that we consider, is that the dark sector may have undergone a strong first-order thermal phase transition in the early universe, possibly producing gravitational waves [37, 38, 39, 40, 41, 42]. Given that in the past decade the latter have finally entered the arena of experimental science [43], it is now particularly timely and important to understand the features of the gravitational-wave spectrum that new-physics models could produce. To this purpose, it should be remarked that, in addition to those produced by non-equilibrium processes at a discontinuous transition, any plasma in thermal equilibrium produces gravitational waves, with an emission rate proportional to the shear viscosity [44]; while the latter can become relatively large at weak coupling, this contribution to the gravitational-wave spectrum remains subdominant with respect to those generated at a first-order phase transition. Gravitational waves produced in a dark sector contribute to the total energy density (and to the effective number of massless degrees of freedom), so the experimental detection of gravitational waves with a characteristic spectrum could provide useful hints about new physics. For example, a first-order phase transition at temperatures of the order of hundreds $\mathrm{GeV}$, i.e., the scale characteristic of electroweak physics, could generate gravitational waves with peak frequencies of the order of $10^{-3}$ Hz, which are potentially within the reach of space-based gravitational-wave observatories [45, 46, 47, 48, 49, 50, 51]. Given that the SM predicts only a smooth crossover at the electroweak scale [52, 53, 54, 55, 56, 57, 58, 59], the experimental detection of gravitational waves in this frequency range could be interpreted as evidence of physics beyond the Standard Model.
@@
Models featuring dark matter candidates in the form of glueball states have been studied in several recent articles; a partial list of references includes refs. [73, 74, 75, 76, 77, 78, 79, 80, 72, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. This family of models can be thought of as examples of the “hidden valley” paradigm [109, 110, 111] (which admits realizations in various theoretical scenarios [112, 113, 114, 115, 116, 117]) and offer phenomenologically attractive features; in particular, while they evade current bounds from colliders [118], they could be coupled to SM particles through portal interactions [119, 120, 121, 122].
The spectrum and related properties of glueballs in $\operatorname{\mathrm{SU}}(N)$ gauge theories have been studied in the lattice regularization in various works [123, 124, 125, 126, 127, 128, 129, 130, 131, 23, 132, 133], but the literature on analogous computations in $\operatorname{\mathrm{Sp}}(N)$ gauge theories is more limited. In fact, with the exception of the smallest symplectic group $\operatorname{\mathrm{Sp}}(1)$ (which is nothing but $\operatorname{\mathrm{SU}}(2)$, and gauge theories based on this group – with or without dynamical matter fields – have already been studied very extensively in lattice simulations, both for the qualitative similarities with QCD and for the potential relevance for physics beyond the SM: for a recent review, see ref. [134, subsection 5.2]), the first lattice study of symplectic gauge theories was reported in ref. [35]; during the past few years, research on this subject has gained significant momentum [135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156].
@@
The structure of this article is as follows. After discussing the formulation of $\operatorname{\mathrm{Sp}}(N)$ gauge theories in the lattice regularization and the method to compute the equation of state in section 2, we present our results for the equilibrium thermodynamic properties of $\operatorname{\mathrm{Sp}}(2)$ theory in section 3, while in section 4 we discuss the possible interpretation of these results in terms of a model for dark matter, and their implications (including, in particular, the possible generation of primordial gravitational waves at the first-order thermal transition of the theory), before recapitulating our findings in section 5. Some general properties of symplectic groups are discussed in appendix A, while appendix B is devoted to technical details about our Markov-chain Monte Carlo algorithm. Throughout this paper we assume natural units, whereby the speed of light in vacuum, the reduced Planck’s constant and Boltzmann’s constant are set to unity.
@@ 2 Formulation of the model
Our non-perturbative study of the equation of state of $\operatorname{\mathrm{Sp}}(2)$ Yang-Mills theory is carried out through a regularization of the theory on a Euclidean, isotropic, hypercubic lattice $\Lambda$ of spacing $a_{\mbox{\tiny{lat}}}$. The physical extent of the system in each of the three spatial directions (labeled by indices $1$, 22 and $3$) is denoted as $L=a_{\mbox{\tiny{lat}}}N_{s}$, while the size of the Euclidean-time direction (labeled by the index 0) is denoted as $a_{\mbox{\tiny{lat}}}N_{\tau}$; furthermore, let $V=L^{3}$ denote the spatial volume of the system. Periodic boundary conditions are imposed along the direction of the four main axes, so that the physical temperature $T$ of the system is given by $T=1/(a_{\mbox{\tiny{lat}}}N_{\tau})$. The simulations at finite temperature are performed on lattices with $N_{s}/N_{\tau}\gtrsim 4$ to suppress finite-volume corrections, while $N_{\tau}=N_{s}$ for the simulations at $T\simeq 0$. For a theory based on a generic $\operatorname{\mathrm{Sp}}(N)$ gauge group, the fundamental degrees of freedom of the theory regularized on the lattice are a set of $4N_{s}^{3}N_{\tau}$ square matrices in the defining representation of the $\operatorname{\mathrm{Sp}}(N)$ group; these matrices have size $2N$ and are defined on the oriented bonds between nearest-neighboring lattice sites; in particular, we denote the matrix on the link from the site $x$ to the site $x+a_{\mbox{\tiny{lat}}}\hat{\mu}$ as $U_{\mu}(x)$. The Euclidean action of the lattice theory is taken to be the Wilson action [36]
@@
In order to express the results extracted from lattice simulations (which are dimensionless numbers, i.e., are measured in units of the appropriate power of the lattice spacing) as physical quantities, one has to set the scale, namely to establish the correspondence between the values of the bare coupling of the lattice theory $g_{0}$ (or, equivalently, the $\beta$ parameter) and the lattice spacing. In this work we set the scale non-perturbatively, by determining the critical $\beta$ at which the deconfinement transition occurs for different values of $N_{\tau}$, interpreting the corresponding values of $1/(a_{\mbox{\tiny{lat}}}N_{\tau})$ as the deconfinement temperature $T_{\mbox{\tiny{c}}}$, and interpolating the numerical results obtained from this procedure to express the lattice spacing as a continuous function of $\beta$ in the range of values that were simulated. [^2] Since the physics of purely gluonic non-Abelian gauge theories in four spacetime dimensions (and without a topological $\theta$ term) depends only on one parameter, setting the scale using a different dimensionful quantity would lead to the same $a_{\mbox{\tiny{lat}}}(\beta)$ function, up to small corrections due to discretization effects. In turn, the determination of the $\beta$ value corresponding to the critical deconfinement temperature is carried out by studying the fourth-order Binder cumulant [175]
@@
[…]
@@
The critical $\beta$ value in the thermodynamic limit is determined through a finite-size scaling analysis, from the crossing of the curves $B_{4}(\beta)$ obtained at different values of the linear spatial size of the system $L$. An alternative quantity that can be used to identify the $\beta$ value corresponding to deconfinement is the susceptibility associated with the spatially-averaged Polyakov loop: [^3] Note that the definition of $\chi_{\mathcal{L}}$ is actually based on the absolute value of $\mathcal{L}$ (rather than $\mathcal{L}$ itself) because all simulations are carried out in finite volume and at finite lattice spacing, and thus involve a strictly finite number of degrees of freedom; as a consequence, any non-vanishing expectation value of $\mathcal{L}$ would simply be an artifact of the finiteness of the sample of configurations. This problem is solved by using $|\mathcal{L}|$, instead of $\mathcal{L}$, in the definition of $\chi_{\mathcal{L}}$.
@@
$\frac{p}{T^{4}}=6N_{\tau}^{4}\int_{\beta_{0}}^{\beta}{\rm{d}}\beta^{\prime}\left(\langle P\rangle_{T}-\langle P\rangle_{0}\right),$ (2.13)
@@
$\int_{x_{0}}^{x_{n}}f(x){\rm{d}}x=h\biggl\{$ $\frac{17}{48}\left[f(x_{0})+f(x_{n})\right]+\frac{59}{48}\left[f(x_{1})+f(x_{n-1})\right]+\frac{43}{48}\left[f(x_{2})+f(x_{n-2})\right]$
@@
Note that, from the pressure and the energy density, it is also possible to derive an estimate for the square of the speed of sound $c_{\mbox{\tiny{s}}}$ in terms of the stiffness:
@@
$L_{\mbox{\tiny{h}}}=\lim_{V\to\infty}\frac{6}{a_{\mbox{\tiny{lat}}}^{4}}\frac{\partial\beta}{\partial\ln a_{\mbox{\tiny{lat}}}}\left(\langle P\rangle_{T_{\mbox{\tiny{c}}}^{-}}-\langle P\rangle_{T_{\mbox{\tiny{c}}}^{+}}\right).$ (2.19)
@@
Finally, before presenting our original results for the thermodynamics of the $\operatorname{\mathrm{Sp}}(2)$ theory, we review the classification of different glueball states whose masses can be computed on the lattice; as will be explained in section 3, this is relevant for our discussion of the equation of state of the theory in the confining phase, which can be modelled in terms of a gas of massive, and approximately non-interacting, glueballs. On a hypercubic lattice, the symmetry under spatial rotations is broken down to rotations by integer multiples of $\pi/2$ around the three main spatial axes, so that glueball states are actually classified according to the irreducible representations of the orientation-preserving octahedral group (the symmetry group of the cube, with $24$ elements falling into five distinct conjugacy classes), which is isomorphic to the group of permutations of four elements $S_{4}$. The latter has five conjugacy classes (respectively containing the identity, the six permutations that can be written as a single cycle of length two, the three permutations that can be written as the product of two disjoint cycles of length two, the eight permutations expressed by a cycle of length three, and the six permutations expressed by a cycle of length four) and hence five irreducible representations; Burnside’s theorem fixes their dimensions to be $1$, 11, 22, 33, and $3$: they are the trivial representation, the sign representation, the two-dimensional irreducible representation, the standard representation, and the product of the standard and sign representations. When regarded as irreducible representations of the orientation-preserving octahedral group, they are usually denoted as $A_{1}$, $A_{2}$, $E$, $T_{1}$, and $T_{2}$, respectively. The restriction of irreducible representations of the $\operatorname{\mathrm{SO}}(3)$ group to the orientation-preserving octahedral group, i.e., the subduced representations, can be decomposed into tensor sums of the various irreducible representations $A_{1}$, $A_{2}$, $E$, $T_{1}$, and $T_{2}$ (for a discussion, see, e.g., refs. [179, 129]). In particular, this implies that the $2J+1$ states of a spin-$J$ representation in the continuum get split among different irreducible representations of the orientation-preserving octahedral group; for example: